# Maximum Length RLL Sequences in de Bruijn Graph

### v1

@@ 6 Conclusion and Discussion
Motivated by an application for space-free quantum key distribution a system based on a simple run-length limited sequences in the de Bruijn graphs is proposed. The maximum length of such sequences is shown to be associated with the number of constrained necklaces. An efficient algorithm to generate a large set of such sequences is proposed and some enumerations related to the length of a maximum length sequences are discussed. Another algorithm that can find efficiently the position of each word of length $n$ in the sequence is also presented.
The exposition of the previous sections can be generalized in a few directions. First, we can consider a non-binary alphabet with the same constraint on the number of zeros, i.e., the longest run of zeros has length at most $s$. The maximum length sequence will be of length which is the total number of words in the necklaces that do not contain a cyclic run of zeros whose length is larger than $s$. A general construction for a large set of maximum length $(n,s)$-sequences can be done similarly as was done in Section 3 using a generalization as done in [7]. Also a maximum length $(n,s)$-sequence in which the position of each word of length $n$ can be efficiently computed is done similarly to the binary case [19].
A different direction is to add more constraints on the sequences. For example, in addition to the requirement that each window of length $n$ has a run of at most $s$ consecutive zeros, we can require that the each window of length $n$ will contain at least $w$ ones.
An $(n,w,s)$-word is an $(n,s)$-word whose weight at least $w$. An $(n,w,s)$-necklace is a necklace for which all its vertices are $(n,w,s)$-words. Let $G_{n}(w,s)$ be the subgraph of $G_{n}$ induced by all the vertices of $G_{n}$ whose representations are $(n,w,s)$-words. The following theorem is proved with the same steps used in Corollary 2.12.
Theorem 6.1.
The length of the circuit ${\cal C}$ in $G_{n-1}(w,s)$ is at most the number of edges in all the $(n,w,s)$-necklaces.
We can merge all the $(n,w,s)$-necklaces with an algorithm which is similar to the algorithm for merging the $(n,s)$-necklaces with adaptation for the threshold on the weight of the windows of length $n$. Theorem 3.1 holds also for this algorithm.
We can also design a more efficient algorithm using a set of words on which the merging of the cycles will be done similarly as was done in the algorithm to merge all the $(n,s)$-necklaces. The ideas will be similar to the ones in [7] with the modification and adaption as was done in Section 3.
The constructed sequences have the potential to be used in other modern applications. A large set of sequences where each sequence can be encoded and decoded efficiently is a major task for the future. Moreover, these algorithms should be also used where the set of constrained words of length $n$ are mapped to a set of information words with a smaller length $k$.

### v2

@@ 6 Conclusion and Discussion
Motivated by an application for space-free quantum key distribution a system based on a simple run-length limited sequences in the de Bruijn graphs is proposed. The maximum length of such sequences is shown to be associated with the number of constrained necklaces. An efficient algorithm to generate a large set of such sequences is proposed and some enumerations related to the length of a maximum length sequences are discussed. Known algorithms to generate one such sequence efficiently are mentioned. Generalizations for larger alphabet or for sequences in which each window of length $n$ has a constrained weight can be easily derived from our exposition.